OpenStudy (anonymous):

Please help me. Algebra 2; 1. If f(x) = 8 – 10x and g(x) = 5x + 4, what is the value of (fg)(–2)? a) –196 b) –168 c) 22 d) 78 2. If f(x) is an odd function and the graph of f(x) includes points in Quadrant IV, which statement about the graph of f(x) must be true? a) It includes points in Quadrant I. b) It includes points in Quadrant II. c) It does not include points in Quadrant I. d) It does not include points in Quadrant II. 3. If f(x) = 3x2 + 1 and g(x) = 1 – x, what is the value of (f – g)(2)? a) 12 b) 14 c) 36 d) 38

OpenStudy (anonymous):

then (fg)=(8-10x)(5x+4) =>(fg)=40x+32-50x^2-40x =>(fg)=32-50x^2 =>(fg)(-2)=32-50(-2)^2 =>(fg)(-2)=32-50*4 =>(fg)(-2)= -168"

OpenStudy (anonymous):

(fg)=(8-10x)(5x+4) (fg)=40x+32-50x^2-40x (fg)=32-50x^2 (fg)(-2)=32-50(-2)^2 (fg)(-2)=32-50*4 (fg)(-2)= -168"

OpenStudy (anonymous):

Thank you both!